After a number of complaints about its directory assistance, a telephone company examined samples of calls to determine the frequency of wrong numbers given to callers. Each of the 8 samples consisted of 131 calls.

6 5 7 10
5 5 7 5

Calculate p one bar (Keep five decimal places in your answer)
Answer #1

P- bar = Average proportion of wrong numbers given to callers

Total number of calls = 131*8 = 1048

Total number of wrong numbers given to callers = 6+5+7+10+5+5+7+5 = 50

Hence,

p - bar = 50/1048 = 0.04771